OpenStudy (anonymous):

Which is the absolute value function corresponding to the graph with these characteristics? Vertex is at (-2, k) V-shaped graph Opens in the downward direction Coefficient a is 4 Left arm passes through point (-1, -3)

OpenStudy (anonymous):

@marissalovescats

OpenStudy (marissalovescats):

Do you have choices?

OpenStudy (anonymous):

y = 4|x − 2| y = 4|x − 2| + 3 y = -4|x + 2| + 3 y = -4|x + 2| − 3

OpenStudy (anonymous):

@marissalovescats

OpenStudy (marissalovescats):

Well if the vertex has -2 in it we can cross out the first 2

OpenStudy (anonymous):

Okay.

OpenStudy (anonymous):

is it C?

OpenStudy (anonymous):

wait no, D.

OpenStudy (marissalovescats):

And since K is positive it has to be +3 So yes C

OpenStudy (popeyes):

Vertex point (-2, k) a = -4 (negative because if open downwards). y = a |x - h| + k y = -4 |x - (-2) | + k y = -4 | x + 2 | + k Left arm passes through point (-3,-1) -1 = -4 | -3 + 2 | + k -1 = -4 | -1 | + k -1 = -4 + k 3 = k Vertex point (-2,k) - (-2,3) a = -4 (negative because opening downwards) y = -4 |x + 2| + 3 So, the answer is C.